The equivalent conductance of 200ml HCL soln has found to be 2×10^3 times its specific conductance. How much water should be added to the soln so that eq        conductance becomes 5×10^3 its specific conductance?

Grade:12

1 Answers

Vikas TU
7 years ago
Eq. Conductance = 2000K = K*1000/N => N1 = 0.5
For 5000K eq. conductance,
5000K = 1000K/N
N2 = 1/5
Now,
N1V1 = N2V2
0.5*200 = 0.2*V2
V2 = 100/0.2 => 500 ml.